A boy throws a ball of mass 0.293 - kg straight upward with an initial speed of 23m/s. When the ball returns to the boy, its speed is 18.24 m/s. How much work does air resistance do on the ball during its flight?


1
Expert's answer
2022-03-16T09:41:45-0400

The work done is equal to change in kinetic energy of the system


"W=\Delta E_k"

"W=\frac{m v_2^2}{2}-\frac{m v_1^2}{2}"

"W=\frac{0.293*18.24^2}{2}-\frac{0.293*23^2}{2}=-28.8\:\rm J"